# Self Model for Embodied Intelligence: Modeling Full-Body Human Musculoskeletal System and Locomotion Control with Hierarchical Low-Dimensional Representation

### 09 Dec 2023 v1

@@ Abstract
Modeling and control of the human musculoskeletal system is important for understanding human motion, developing embodied intelligence, and optimizing human-robot interaction systems. However, current open-source models are restricted to a limited range of body parts and often with a reduced number of muscles. There is also a lack of algorithms capable of controlling over 600 muscles to generate reasonable human movements. To fill this gap, we build a comprehensive musculoskeletal model with 90 body segments, 206 joints, and 700 muscle-tendon units, allowing simulation of full-body dynamics and interaction with various devices. We develop a new algorithm using low-dimensional representation and hierarchical deep reinforcement learning to achieve state-of-the-art full-body control. We validate the effectiveness of our model and algorithm in simulations and on real human locomotion data. The musculoskeletal model, along with its control algorithm, will be made available to the research community to promote a deeper understanding of human motion control and better design of interactive robots.
@@ I Introduction
The human musculoskeletal system is a complex dynamic system with high spatio-temporal degrees of freedom. While current technology cannot effectively measure the dynamic characteristics of the human body as a whole, modeling and simulation are very important for the understanding of human motion control and human factors in human-robot interaction. Dynamic modeling could also facilitate humanoid design and provide a human self-model for embodied intelligence. A musculoskeletal model is built from anatomical and biomechanical data of the human body. By simulating human motion with a full-body musculoskeletal model, it is possible to gain insight into the dynamics of the human body and the factors that influence motion, such as muscle activation patterns, joint stability, and limb coordination. Control algorithms for the movements of a human musculoskeletal model will contribute to our understanding of human motion generation and human-robot collaboration. We can use simulation data instead of real experimental data to design and optimize robots, such as humanoids, exoskeletons, prosthetics, and other human-interactive devices.
Much of our daily movement requires the use of the entire body. For example, walking and running utilize swings and stances of the lower limbs to provide propulsion, combined with swings of the upper limbs to balance and accelerate the body, and trunk muscles are needed to maintain stability of the torso. However, there is no open-source musculoskeletal model that represents the entire human body. Moreover, existing algorithms can hardly handle the very high dimensional control problem of full-body human musculoskeletal motion.
The development of a musculoskeletal model of the full human body, and the control of its movement generation is profoundly difficult due to the following challenges. First, building a human musculoskeletal model requires detailed quantitative knowledge of human anatomy and physiology. Second, the muscle control parameter space is complex, where over 600 skeletal muscles control hundreds of joints, leading to high dimensionality and redundancy of the system [1, 2]. Third, the dynamical property of neuro-muscle actuators is non-linear and inconstant[3, 4].
@@
Here, we propose a full-body human musculoskeletal model with 90 body segments, 206 joints and 700 muscle-tendon units for neuromuscular-driven dynamic simulation, as shown in Fig.1. The model can be embedded into physics simulation engines like MuJoCo[5] and OpenSim[6], establishing a framework that can be used to perform neuromusculoskeletal control simulation of human full-body movements.
We also develop a hierarchical deep reinforcement learning algorithm with a low-dimensional representation, Two-Stage Hierarchical Training (TSHT), which is capable of controlling the model to generate biologically plausible movements. We make both the musculoskeletal model and our control algorithm available to the research community with the goal of fostering the development of more accurate models, the design of better assistive robots, and a better understanding of human neuromuscular motion control for embodied intelligence.
Our Contributions. We build a dynamic model of full-body human musculoskeletal system with physiologically feasible wrapping geometry for segments, joints, and multi-muscle-tendon units. We embed the comprehensive model in a simulation environment that supports contact interaction simulations, providing a scalable, adaptable, and computationally efficient simulation framework for data-driven musculoskeletal system control. We propose a deep reinforcement learning algorithm, two-stage hierarchical training, for the control of the high-dimensional and highly nonlinear full-body musculoskeletal model. Our control algorithm can generate desired motions in full-body musculoskeletal model.
@@ II Related Work
Human Musculoskeleton Modeling. Many previous studies focused on the musculoskeletal models actuated by Hill-type muscles[3] for specific human body parts. An upper extremity model has been built and validated[7, 8, 9]. Lower extremity models have been continuously improved and used in the field[10, 11, 12, 13, 14]. Some researchers built the model of the trunk[15] and extended it with simple muscles actuated limbs[16]. Other researchers created models with muscles of the torso and limbs, but the models do not cover a large range of important skeletal muscles and joints of the torso[17, 18]. These models were mainly built on physics-based musculoskeletal simulation platforms. Representative research efforts include OpenSim [6], Anybody[19], etc. These platforms provide researchers with the means to create sophisticated musculoskeletal models. However, running dynamic simulations on them is computationally expensive. Some recently developed physics engines[5, 20, 21, 22] are becoming popular for robotic simulation, and they are capable for efficient dynamic simulations that support contact interactions. MuJoCo’s modeling approach is able to provide a simplified and fully supported interface for muscle modeling among them. Efforts have been made to convert the musculoskeletal models from OpenSim to MuJoCo, resulting in acceleration of simulation speed on the model by two orders of magnitude[23, 24, 25].
@@
Motion Control of the Musculoskeletal System. Researchers have begun to explore the use of biological actuators for control human musculoskeletal systems. Evolutionary algorithms were used to optimize control parameters for a 2D lower limb model comprising 16 muscle-tendon units[26], and previous work [27] developed a more comprehensive musculoskeletal model. However, these methods fell short when handling complex models. Recently, a hierarchical reinforcement learning algorithm, coupled with imitation learning, was applied to a musculoskeletal model featuring 346 muscles[17]. Xu et al.[28] introduced a high-performance differentiable simulator and an algorithm that effectively enhanced control over the human musculoskeletal model by leveraging gradient information. A review was conducted on the application of reinforcement learning in neuromechanical simulations[29]. Beyond the human musculoskeletal model, a 3D musculoskeletal model of an ostrich was constructed using the MuJoCo engine[30]. This model successfully performed various locomotion tasks with deep reinforcement learning algorithm TD4. Schumacher et al.[31] proposed the DEP-RL algorithm to attain superior control over the aforementioned ostrich model. Furthermore, the concept of reward shaping to achieve natural walking in musculoskeletal models of human lower limbs was introduced[32]. Multi-task learning[33] and synergistic action representation[34, 35] have demonstrated their effectiveness in the motion control of specific musculoskeletal body parts.
@@
In the open-source musculoskeletal modeling community, there is no full-body human musculoskeletal model with more than 600 muscles. Control algorithms that can drive the high-dimensional musculoskeletal models to produce realistic motions are still lacking.
@@ III Musculoskeletal System
As demonstrated in Fig.1, our full-body musculoskeletal model includes 90 rigid body segments, 206 joints, and 700 muscle-tendon units. We use musculoskeletal models[8, 9, 12, 14, 15] on OpenSim platform as the references to set anatomical parameters of the following models of body segments, joints and muscle-tendon units, following the pipeline of [23]. We also embed the MARCH exoskeleton model in simulation to demonstrate human-machine interactive walking.
@@ III-A Skeletal Modeling
$M(q)\ddot{q}+c(q,\dot{q})=J_{m}^{T}f_{m}(a)+J_{ext}^{T}\tau_{ext}$ (1)
where $q$ represents the generalized coordinates of joints, $M(q)$ is the mass distribution matrix, $c(q,\dot{q})$ is Coriolis and gravitational forces. $f_{m}(a)$ is the vector representing muscle forces generated by all muscle-tendon units, and $\tau_{ext}$ is the external force from the interaction with environments. $J_{m}$ and $J_{ext}$ are Jacobian matrices that map forces to the space of generalized coordinates.
@@ III-B Muscle Modeling
Muscle-Tendon Pathway. These muscle-tendon units typically originate and are inserted at pathway points attached to different body segments. For wide muscle attachments, it will be modeled as multiple independent muscle-tendon units. Thus, the total number of units, 700, in our model is larger than the common biomedical estimations for the number of skeletal muscles, which is around 640 [37]. Wrapping geometry surfaces are added to the pathways to model physical constraints imposed by bones and soft tissue when needed. Our multi-muscle-tendon and wrapping geometry model is demonstrated in Fig. 2(a) and Fig. 2(b).
@@
Neuro-Muscle Dynamics. The activation-contraction dynamics of muscles exhibit non-linearity and temporal delay, thereby posing challenges to neuromuscular control. With the employment of the Hill-type muscle model[3], the maximum isometric force, the tendon length, and the range of muscle length of each unit is defined in the model. The force produced by a single muscle-tendon unit can be formulated as
$f_{m}(a)=f_{max}\cdot[F_{l}(l_{m})\cdot F_{v}(v_{m})\cdot a+F_{p}(l_{m})],$ (2)
where $f_{max}$ stands for the maximum isometric muscle force and $a,l_{m},v_{m}$ respectively stand for the activation, normalized length and normalized velocity of the muscle. $F_{l}$ and $F_{v}$ represent force-length and force-velocity functions fitted using data from biomechanical experiments[38]. The muscle activation $a$ is calculated by a first-order nonlinear filter in simulation as Eq. (3) where $u$ is the input control signal of the musculoskeletal model. $(\tau_{act},\tau_{deact})$ is a time constant for activate and deactivate latency with defaults $(10ms,40ms)$.
$\frac{\partial a}{\partial t}=\frac{u-a}{\tau(u,a)},\tau(u,a)=\begin{cases}\tau_{act}(0.5+1.5a)&u>a\\ \frac{\tau_{deact}}{0.5+1.5a}&u\leq a\end{cases}$ (3)
@@ IV-A Problem formulation for Musculoskeletal Model Control
We formulate the control problem of the musculoskeletal model as a Markov Decision Processes, denoted by $\langle\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\rho_{0},\gamma\rangle$, where $\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the continuous space of all valid states, $\mathcal{A}\subseteq\mathbb{R}^{m}$ represents the continuous space of all valid actions, where $\mathcal{R}$ stands for the reward function, where $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. $\mathcal{P}$ stands for the state transfer probability function, where $\mathcal{P}(s^{\prime}|s,a)$ denotes the probability of an agent taking an action $a$ to transfer from the current state $s$ to the state $s^{\prime}$, $\rho_{0}$ is the probability distribution of the initial state with $\sum_{s\in\mathcal{S}}\rho_{0}(s)=1$ and $\gamma\in[0,1)$ is the discount factor. We adopt an off-policy RL algorithm, Soft Actor-Critic (SAC)[39], to optimize our policy. This algorithm aims to obtain the optimal policy by maximizing the accumulative discounted return, $\sum_{t}\gamma^{t}R_{t}$ and the policy entropy, $\mathcal{H}(\pi)$, where
@@
State space $\mathcal{S}$ consists of $\{t,q,\dot{q},f_{m},l_{m},\dot{l}_{m},act,\mathcal{E}\}$, represents simulation time, joint position, joint velocity, muscle force, muscle length, muscle velocity, muscle activation, and other task-specific observations respectively. Task-specific observations will be introduced in the experiments section. Action space $\mathcal{A}$ consists of muscle  excitation $u$. We can also call muscle excitation as motor neuron signals.
@@ IV-B Motion Imitation
The reference generator shown in Fig.4 can generate each considered joint position $q^{r}$ and velocity $\dot{q}^{r}$ taking motion capture data from different tasks. The optical motion capture system calculated the position of markers, based on which we perform an inverse kinematic to reconstruct motions. We then obtain a time series of joint angles. Taking as input the distance between the actual joint state $\left\{q,\dot{q}\right\}$ and the desired joint state $\left\{q^{r},\dot{q}^{r}\right\}$, the policy is able to accomplish all considered tasks efficiently.
@@ IV-C Hierarchical Action Space Representation
As a result of the high dimensional nature of the action space, reinforcement learning algorithms in previous studies[30, 31] failed to retain control of our full-body musculoskeletal models in the generation of locomotion. As a result, the insufficiently controlled model appears to fall easily during simulation. Given the vastness of the action space, the key to resolving this problem lies on the representation of the action space. We use the physiological muscle synergy and low-dimensional representation of muscle activation to speed up the policy learning process.
@@ V-A Experimental Setup
In this task, the reward consists of three parts, where $w_{q}$, $w_{\dot{q}}$ and $w_{c}$ are are the weights of the reward for position tracking, velocity tracking and concerned joint position tracking respectively. $w_{q}=1,w_{\dot{q}}=5e^{-3},w_{c}=1$. The musculoskeletal model may be initialized at any point throughout a trajectory cycle. The task does not have task-specific observation.
Human-Exo interaction. We also design a task on human interaction with the MARCH exoskeleton. In this experiment, we expected the model to walk in the exoskeleton with minimal contact force in legs. Similarly, we use trajectory tracking for its upper limbs. The joint angles of the exoskeleton are tracking the reference trajectory of lower limbs by a PID controller. And we use spring elements with high stiffness to model the bandage interaction between legs and exoskeleton.
@@
In this task, the reward comprises a penalty for contact force and a reward for tracking the upper limb trajectory, where $w_{m}$ and $w_{c}$ are the weights of the corresponding rewards. $w_{m}=10,w_{c}=5e^{-2}$. The task-specific observations for this task include the distance between the ends of the elastic elements on the leg and the exoskeleton.

### 09 Mar 2024 v2

@@ Abstract
Modeling and control of the human musculoskeletal system is important for understanding human motor functions, developing embodied intelligence, and optimizing human-robot interaction systems. However, current open-source models are restricted to a limited range of body parts and often with a reduced number of muscles. There is also a lack of algorithms capable of controlling over 600 muscles to generate reasonable human movements. To fill this gap, we build a musculoskeletal model with 90 body segments, 206 joints, and 700 muscle-tendon units, allowing simulation of full-body dynamics and interaction with various devices. We develop a new algorithm using low-dimensional representation and hierarchical deep reinforcement learning to achieve state-of-the-art full-body control. We validate the effectiveness of our model and algorithm in simulations with real human locomotion data. The musculoskeletal model, along with its control algorithm, will be made available to the research community to promote a deeper understanding of human motion control and better design of interactive robots.
@@ I Introduction
The human musculoskeletal system is a complex dynamic system with high spatio-temporal degrees of freedom. While current technological advancements cannot effectively measure the dynamic characteristics of the human body as a whole, modeling and simulation are very important for the understanding of human motion control and human factors in human-robot interaction. Dynamic modeling can also facilitate the design of humanoid robots and provide a self-model of human for embodied intelligence. A musculoskeletal model is built from anatomical and biomechanical data of the human body. By simulating human motion with a musculoskeletal model, it is possible to gain insight into the dynamics of the human body and the factors that influence motion, such as muscle activation patterns, joint stability, and limb coordination. We can use simulation data instead of hard-to-obtain real data to design and optimize robots, such as humanoids, exoskeletons, prosthetics, and other human-interactive devices.
Many of our daily movements require the use of the entire body. For instance, walking and running involve the coordinated movement of the lower limbs for propulsion, complemented by the swinging motion of the upper limbs for balance, while the trunk muscles contribute to torso stability. However, there is no open-source musculoskeletal model that represents the entire human body. Constructing a full-body human musculoskeletal model enables comprehensive characterization of the action space and state space in human movement simulations. The dynamics of the model improves our understanding of human motor intelligence. Furthermore, current algorithms struggle to address the highly dimensional control problem associated with full-body human musculoskeletal motion.
The modeling and control of the musculoskeletal model of full human body is profoundly difficult due to the following challenges. First, building a human musculoskeletal model requires detailed quantitative knowledge of anatomy and physiology. Second, the muscle control parameter space is complex, where over 600 skeletal muscles control hundreds of joints, leading to high dimensionality and redundancy of the system[1, 2]. Third, the dynamical property of neuro-muscle actuators is non-linear and inconstant[3, 4].
@@
Here, we propose a full-body human musculoskeletal model with 90 body segments, 206 joints and 700 muscle-tendon units for neuromuscular-driven dynamic simulation, as shown in Fig.1. The model can be imported into open-source physics simulation engines like MuJoCo[5] and OpenSim[6], establishing a framework that can be used to perform neuro-musculo-skeletal control simulation of human full-body movements.
We also develop a hierarchical deep reinforcement learning (RL) algorithm with a low-dimensional representation, Two-Stage Hierarchical Training (TSHT), which is capable of controlling the model to generate biologically plausible movements. We make both the musculoskeletal model and our control algorithm available to the research community with the goal of fostering the development of more accurate models, the design of better assistive robots, and a better understanding of human neuromuscular motion control for embodied intelligence.
Our Contributions. We build a dynamic model of full-body human musculoskeletal system with physiologically feasible body segments, joints, and multi-muscle-tendon units. We embed the full-body model in simulation environments that support contact interaction simulations, providing a scalable, adaptable, and computationally efficient simulation framework for data-driven musculoskeletal system control. We propose a deep reinforcement learning algorithm, two-stage hierarchical training, for the control of the high-dimensional and highly nonlinear full-body musculoskeletal model. Our control algorithm can generate desired motions in full-body musculoskeletal model.
@@ II Related Work
Human Musculoskeleton Modeling. Many previous studies focused on the musculoskeletal models actuated by Hill-type muscles[3] for specific human body parts. An upper extremity model has been built and validated[7, 8, 9]. Lower extremity models have been continuously improved and used in the field[10, 11, 12, 13, 14]. Some researchers built the model of the trunk[15] and extended it with simple muscles-actuated limbs[16]. Other researchers created models with muscles of the torso and limbs, but the models do not cover a large range of important skeletal muscles and joints of the torso[17, 18]. These models were mainly built on physics-based musculoskeletal simulation platforms. Representative research efforts include OpenSim [6], Anybody[19], etc. However, running dynamic simulations on them is computationally expensive. Some recently developed physics engines[5, 20, 21, 22] are becoming popular for robotic simulation, and they are capable for efficient dynamic simulations that support contact interactions. MuJoCo’s modeling approach is able to provide a simplified and fully supported interface for muscle modeling among them. Efforts have been made to convert the musculoskeletal models from OpenSim to MuJoCo, resulting in acceleration of simulation speed on the model by two orders of magnitude[23, 24, 25].
@@
Motion Control of the Musculoskeletal System. Researchers have explored the use of biological actuators for controlling human musculoskeletal systems. Evolutionary algorithms were used to optimize control parameters for a 2D lower limb model comprising 16 muscle-tendon units[26], and previous work [27] developed a more complex musculoskeletal model. However, these methods fell short when handling complex models. Recently, a hierarchical RL algorithm, coupled with imitation learning, was applied to a musculoskeletal model featuring 346 muscles[17]. Xu et al.[28] introduced a high-performance differentiable simulator and an algorithm that effectively enhanced control over the human musculoskeletal model by leveraging gradient information. A review was conducted on the application of RL in neuromechanical simulations[29]. Beyond the human musculoskeletal model, a 3D musculoskeletal model of an ostrich was constructed using the MuJoCo engine[30]. This model successfully performed various locomotion tasks with deep RL algorithm TD4. Schumacher et al.[31] proposed the DEP-RL algorithm to attain superior control over the aforementioned ostrich model. Furthermore, the concept of reward shaping to achieve natural walking in musculoskeletal models of human lower limbs was introduced[32]. Multi-task learning[33] and synergistic action representation[34, 35] have demonstrated their effectiveness in the motion control of specific musculoskeletal body parts.
@@
In the open-source musculoskeletal modeling community, the current model with the largest number of muscles contains 552 individual muscle units[16], which lacks the coverage of limb muscles to complete the simulation of full-body movements. Control algorithms that can drive the high-dimensional musculoskeletal models to produce realistic motions are still lacking.
@@ III Musculoskeletal System
As demonstrated in Fig.1, our full-body musculoskeletal model includes 90 rigid body segments, 206 joints, and 700 muscle-tendon units. The model can simulate the joint movements of the torso and limbs driven by skeletal muscles. We use musculoskeletal models[8, 9, 12, 14, 15] on OpenSim platform as the references to set anatomical parameters of the following models of body segments, joints and muscle-tendon units, following the pipeline of [23]. We also embed the MARCH exoskeleton model in simulation to demonstrate human-machine interactive walking.
@@ III-A Skeletal Modeling
$M(q)\ddot{q}+c(q,\dot{q})=J_{m}^{T}f_{m}(act)+J_{c}^{T}f_{c}+\tau_{ext},$ (1)
where $q$ represents the generalized coordinates of joints, $M(q)$ is the mass distribution matrix, $c(q,\dot{q})$ is Coriolis and gravitational forces. $f_{m}(act)$ is the vector representing muscle forces generated by all muscle-tendon units, and is determined by muscle activations ($act$). $f_{c}$ is the constraint force, and $\tau_{ext}$ is the external torque from the interaction with environments. $J_{m}$ and $J_{c}$ are Jacobian matrices that map forces to the space of generalized coordinates.
@@ III-B Muscle Modeling
Muscle-Tendon Pathway. These muscle-tendon units typically originate and are inserted at pathway points attached to different body segments. For wide muscle attachments, it will be modeled as multiple independent muscle-tendon units. Thus, the total number of units, 700, in our model is larger than the common biomedical estimations for the number of skeletal muscles, which is around 640 [37]. Wrapping geometry surfaces are added to the pathways to model physical constraints imposed by bones and soft tissue when needed. Our multi-muscle-tendon and wrapping geometry models are demonstrated in Fig. 2(a) and Fig. 2(b).
@@
Neuro-Muscle Dynamics. The activation-contraction dynamics of muscles exhibit non-linearity and temporal delay, thereby posing challenges to neuromuscular control. With the employment of the Hill-type muscle model[3], the maximum isometric force, the tendon length, and the optimal fiber length of each unit is defined in the model. The force produced by a single muscle-tendon unit can be formulated as
$f_{m}(act)=f_{max}\cdot[F_{l}(l_{m})\cdot F_{v}(v_{m})\cdot act+F_{p}(l_{m})],$ (2)
where $f_{max}$ stands for the maximum isometric muscle force and $act,l_{m},v_{m}$ respectively stand for the activation, normalized length and normalized velocity of the muscle. $F_{l}$ and $F_{v}$ represent force-length and force-velocity functions fitted using data from biomechanical experiments[38]. The muscle activation $act$ is calculated by a first-order nonlinear filter in simulation as Eq. (3) where $u$ is the input control signal of the musculoskeletal model. $(\tau_{\text{act}},\tau_{\text{deact}})$ is a time constant for activate and deactivate latency with defaults $(10ms,40ms)$.
$\frac{\partial act}{\partial t}=\frac{u-act}{\tau(u,act)},\tau(u,act)=\begin{cases}\tau_{\text{act}}(0.5+1.5act)&u>act\\ \frac{\tau_{\text{deact}}}{0.5+1.5act}&u\leq act\end{cases}$ (3)
@@ IV-A Problem formulation for Musculoskeletal Model Control
We formulate the control problem of the musculoskeletal model as a Markov Decision Processes, denoted by $\langle\mathcal{S},\mathcal{A},r,p,\rho_{0},\gamma\rangle$, where $\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the continuous space of all valid states, $\mathcal{A}\subseteq\mathbb{R}^{m}$ represents the continuous space of all valid actions. $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward funtion. $p$ stands for the state transfer probability function, where $p(s^{\prime}|s,a)$ denotes the probability of an agent taking an action $a$ to transfer from the current state $s$ to the state $s^{\prime}$, $\rho_{0}$ is the probability distribution of the initial state with $\sum_{s\in\mathcal{S}}\rho_{0}(s)=1$ and $\gamma\in[0,1)$ is the discount factor. We adopt an off-policy RL algorithm, Soft Actor-Critic (SAC)[39], to optimize our policy. This algorithm aims to obtain the optimal policy by maximizing the accumulative discounted return, $\sum_{t}\gamma^{t}R_{t}$ and the policy entropy, $\mathcal{H}(\pi)$, where
@@
State space $\mathcal{S}$ consists of $\{t,q,\dot{q},f_{m},l_{m},\dot{l}_{m},act,\mathcal{E}\}$, represents simulation time, joint position, joint velocity, muscle force, muscle length, muscle velocity, muscle activation, and other task-specific observations respectively. Task-specific observations are introduced in section V-A. Action space $\mathcal{A}$ consists of muscle excitation $u$. We can also call muscle excitation as motor neuron signals.
@@ IV-B Motion Imitation
The reference generator shown in Fig. 4 can generate reference joint position $q^{r}$ and velocity $\dot{q}^{r}$ taking motion capture data from different tasks. The optical motion capture system calculated the position of markers, based on which we perform an inverse kinematic to reconstruct motions. We then obtain a time series of joint angles. Taking the distance between the actual joint state $\left\{q,\dot{q}\right\}$ and the desired joint state $\left\{q^{r},\dot{q}^{r}\right\}$ as input, the policy is able to accomplish all considered tasks efficiently.
@@ IV-C Hierarchical Action Space Representation
Due to the high-dimensional nature of the action space, previous algorithms[30, 31] struggled to maintain control over our full-body musculoskeletal model during locomotion generation. Consequently, the inadequately controlled models frequently collapsed during simulation. Given the vastness of the action space, resolving this issue hinges on the representation of the action space. We employ physiological muscle synergy and a low-dimensional representation of muscle activation to expedite the policy learning process.
@@ V-A Experimental Setup
The reward consists of three parts, where $w_{q}$, $w_{\dot{q}}$ and $w_{c}$ are the weights of the reward for position tracking, velocity tracking and concerned joint position tracking respectively. $w_{q}=1,w_{\dot{q}}=5e^{-3},w_{c}=1$. The musculoskeletal model may be initialized at any point throughout a trajectory cycle. The task does not have task-specific observation.
Human-Exo interaction. We also design a task on human interaction with the MARCH exoskeleton. In this environment, we expected the model to walk in the exoskeleton with minimal contact force in legs. Similarly, we use trajectory tracking for its upper limbs. The joint angles of the exoskeleton are tracking the reference trajectory by a PID controller. We use spring elements with high stiffness to model the bandage interaction between legs and exoskeleton.
@@
The reward comprises a penalty for contact force and a reward for tracking the upper limb trajectory, where $w_{m}$ and $w_{c}$ are the weights of the corresponding rewards. $w_{m}=10,w_{c}=5e^{-2}$. The task-specific observations for this task include the distance between the ends of the elastic elements on the leg and the exoskeleton.